OpenStudy (anonymous):

Find the measure of the given angles. An angle whose supplement is 3 times the measure of it's complement.

OpenStudy (anonymous):

(180-X)=3(90-X) 180-X = 270-3X 180-270 = -3X+X - 90 = -2X X = -90/-2 X = 46 degree If I deserve medal then give me one.

OpenStudy (anonymous):

wait I miss 180=3(90-X) 180 = 270-3X 180-270 = -3X - 90 = -3X X = -90/-3 X = 30 degree

OpenStudy (anonymous):

T Y

OpenStudy (anonymous):

thx